OpenStudy (anonymous):

the front row of a stadium has 25 seats. each of the other rows has two more seats than the row in front of it. how many seats are there altogether in the first 20 rows?

OpenStudy (anonymous):

25+(x*2)

OpenStudy (anonymous):

oh wow, thanks for the quick response!

OpenStudy (anonymous):

not finished

OpenStudy (anonymous):

@timo86m :can u explain how did u get that ?

OpenStudy (anonymous):

im studying for the SATs and am struggling greatly with permutations

OpenStudy (anonymous):

sequences*

OpenStudy (anonymous):

@Megichu : U know A.Sequence basics ?

OpenStudy (anonymous):

2+2+2+2...=19*2 =38 25+(19*25)+38

OpenStudy (anonymous):

i understand what the basic formula is, but interpreting the problem at times can be difficult

OpenStudy (anonymous):